Evaluate the integral of (5e)^(3x) dx


1
Expert's answer
2021-03-30T07:36:41-0400

∫(5e)3x dx = ∫(eln5 *e)3x dx = ∫(e1+ln5)3x dx = ∫e(1+ln5)3x dx = e3(1+ln5)x / [3(1+ln5)] + C= e3x e3xln5 / [3(1+ln5)] = e3x 53x / [3(1+ln5)] + C

Answer

∫(5e)3x dx = e3x 53x / [3(1+ln5)] + C